OpenStudy (anonymous):

If g(x) = 2x2 + bx + 5 and g(1) = 4, what is the value of g(-1)? 1 2 3 7 10

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle g(x)=2x^2+bx+5 }\) correct?

OpenStudy (anonymous):

yes sir

OpenStudy (solomonzelman):

Your given that \(\large\color{red}{ \color{black}{g(}1\color{black}{)=}\color{blue}{4} }\), so lets apply this. \(\large\color{black}{ \displaystyle g(x)=2x^2+bx+5 }\) for any function g(x)=function with x's. when you do g(1), every single x is replaced by a 1. This gets us: \(\large\color{black}{ \displaystyle g(\color{red}{1})=2(\color{red}{1})^2+b(\color{red}{1})+5 }\)

OpenStudy (solomonzelman):

Now, we know that \(\large\color{red}{ \color{black}{g(}1\color{black}{)=}\color{blue}{4} }\) So, we can substitute 4 instead of g(1), the following way: \(\large\color{black}{ \displaystyle \color{blue}{4}=2(\color{red}{1})^2+b(\color{red}{1})+5 }\)

OpenStudy (solomonzelman):

All you need to do know is to simplify and solve for b:)

OpenStudy (anonymous):

i still need help sir solving for b @SolomonZelman

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \color{blue}{4}=2(\color{red}{1})^2+b(\color{red}{1})+5 }\) ok, what is `2(1)²` equal to?

OpenStudy (anonymous):

2?

OpenStudy (solomonzelman):

yes

OpenStudy (solomonzelman):

So lets write that \(\large\color{black}{ \displaystyle \color{blue}{4}=2+b(\color{red}{1})+5 }\)

OpenStudy (solomonzelman):

b(1) is b•1, and that is just b. Right?

OpenStudy (anonymous):

yes

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \color{blue}{4}=2+b+5 }\)

OpenStudy (solomonzelman):

can you solve this, or need a little more help?

OpenStudy (anonymous):

i need a little more help please

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \color{blue}{4}=2+b+5 }\) \(\large\color{black}{ \displaystyle \color{blue}{4}=7+b }\)

OpenStudy (solomonzelman):

agree?

OpenStudy (anonymous):

but do we not consider the b to be a imaginary 1 ?

OpenStudy (solomonzelman):

lol, it seems as though you are having a bad day right now...

OpenStudy (solomonzelman):

4=7+b ust subtract 7 from both sides

OpenStudy (anonymous):

lol no sir i just want to understand it fully

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle 4=7+b }\) \(\large\color{black}{ \displaystyle 4\color{magenta}{-7}=7+b \color{magenta}{-7} }\) \(\large\color{black}{ \displaystyle 4\color{magenta}{-7}=\cancel{7}+b \cancel{\color{magenta}{-7}} }\)

OpenStudy (solomonzelman):

b=?

OpenStudy (anonymous):

3

OpenStudy (solomonzelman):

not exactly, but close...

OpenStudy (solomonzelman):

b can be negative too:)

OpenStudy (anonymous):

but my answer choices are not negative

OpenStudy (solomonzelman):

then they are wrong, because the answer is -3.

OpenStudy (solomonzelman):

Oh, we are not done yet....

OpenStudy (solomonzelman):

the answer choices are for g(-1)

OpenStudy (solomonzelman):

and -3 that we found just now is the value of b

OpenStudy (anonymous):

yes

OpenStudy (solomonzelman):

Now we know our function entirely. g(x) = 2x² - 3x + 5

OpenStudy (solomonzelman):

And now find g(-1), by plugging -1 instead of x.

OpenStudy (solomonzelman):

g(x) = 2x² - 3x + 5 g(-1) = 2(-1)² - 3(-1) + 5=?

OpenStudy (anonymous):

g(-1)=2 -3+5=4

OpenStudy (anonymous):

???

OpenStudy (anonymous):

or is it 10?

OpenStudy (solomonzelman):

oh wait

OpenStudy (anonymous):

i think it is 10 sir

OpenStudy (solomonzelman):

g(-1) = 2(-1)² - 3(-1) + 5 g(-1) = 2 - - 3 + 5 g(-1) = 2 + 3 + 5 g(-1) = 10

OpenStudy (solomonzelman):

correct

OpenStudy (anonymous):

thank you so much you are awesome :)

OpenStudy (solomonzelman):

You are welcome!